If x + y + z = 0, show that r + y 9 + z3 • 3zyz.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Consider the identity r3 + y3 + z3 - 3xy:z • = (x + y + z) (x2 + y2 + z2 - xi -Yz -.u) •••(i) If x + y + z = 0, then